OpenStudy (ghdoru11):

Find the area of the plane surface bounded by the curves: y=sqrt{10}x and y^2=10x

OpenStudy (dumbcow):

Find where they intersect \[\sqrt{10} x = \sqrt{10} \sqrt{x}\] \[x = \sqrt{x}\] x = 0, 1